OpenStudy (anonymous):

One pipe can fill a tank in 4 hours, and another pipe can fill the tank in 6 hours. How long will it take to fill the tank if both pipes are used ?

OpenStudy (radar):

Use product over sum: 6x4/(6+4)=2.4 hours

OpenStudy (anonymous):

So after 1 hour, the first pipe can fill the tank 1/4 of the way up and the second pipe can fill the tank up 1/6 of the way up. Together, they will fill 1/4+1/6 of the tank up in one hour. That is 5/12 of the tank is filled after one hour. So you want to set a proportion now. 5/12s / 1 hour =1 tank / x hours, cross multiple and solve for x.

OpenStudy (anonymous):

can u show me how to do this problem?

OpenStudy (amistre64):

How much is filled in 1 hour? P1 = 1/4 P2 = 1/6 P1 + P2 = amount filled in one hour. 1/4 + 1/6 = (6+4)/(24) = 5/12 1hr: 5/12 2hr 10/12 with 2/12 left.

OpenStudy (amistre64):

Better yet: (5/12)/60 = 1/144 per minute. a full tank is 144/144 so it takes 144 minutes to fill it. convert back to hours 144/60= 2.4 hrs